# Adaptive Multi-modal Fusion of Spatially Variant Kernel Refinement with Diffusion Model for Blind Image Super-Resolution

### v1

@@ Abstract
Pre-trained diffusion models utilized for image generation encapsulate a substantial reservoir of a priori knowledge pertaining to intricate textures. Harnessing the potential of leveraging this a priori knowledge in the context of image super-resolution presents a compelling avenue. Nonetheless, prevailing diffusion-based methodologies presently overlook the constraints imposed by degradation information on the diffusion process. Furthermore, these methods fail to consider the spatial variability inherent in the estimated blur kernel, stemming from factors such as motion jitter and out-of-focus elements in open-environment scenarios. This oversight results in a notable deviation of the image super-resolution effect from fundamental realities. To address these concerns, we introduce a framework known as Adaptive Multi-modal Fusion of Spatially Variant Kernel Refinement with Diffusion Model for Blind Image Super-Resolution (SSR). Within the SSR framework, we propose a Spatially Variant Kernel Refinement (SVKR) module. SVKR estimates a Depth-Informed Kernel, which takes the depth information into account and is spatially variant. Additionally, SVKR enhance the accuracy of depth information acquired from LR images, allowing for mutual enhancement between the depth map and blur kernel estimates. Finally, we introduce the Adaptive Multi-Modal Fusion (AMF) module to align the information from three modalities: low-resolution images, depth maps, and blur kernels. This alignment can constrain the diffusion model to generate more authentic SR results. Quantitative and qualitative experiments affirm the superiority of our approach, while ablation experiments corroborate the effectiveness of the modules we have proposed.
@@ 1 Introduction
Image super-resolution is a foundational pursuit within the domain of computer vision, focusing on the reconstruction of high-resolution images from their low-resolution counterparts to significantly augment pixel density [7, 6]. This intricate task involves the inference and restoration of high-frequency details that are often obscured within the constraints of limited pixel landscapes. Mathematically, the regression process from a high-resolution image (HR) to a low-resolution image (LR) is modeled as follows:
@@
where $k$ symbolizes the blur kernel, $\otimes$ denotes the convolution operation, $\downarrow_{s}$ signifies a downsampling by a factor of $s$, $n$ represents noise. Over the past decade, methods grounded in CNN or transformer have made remarkable strides [12, 14, 35]. Traditional super-resolution methods assume this degradation is characterized by a specific blur kernel, such as a bicubic kernel [36]. However, when a well-trained SR model recevies LR images away from its training distribution, often suffers significant performance degradation [25]. To tackle this problem, blind super-resolution tasks aim at image super-resolution with unknown degradation processes [29].
As depicted in Figure 2, spatially invariant estimated blur kernels, incorporating a priori degradation information, are employed in the super-resolution process to enhance performance and achieve a finite solution space for the LR-HR image pair [5, 8]. In real-world imaging, factors limiting image resolution extend beyond specific blur kernels to encompass motion blur, depth variation in imaging objects, and optical characteristics of imaging devices [65]. The use of a spatially invariant blur kernel becomes inadequate in accurately describing this multifaceted process, yielding unsatisfactory results post super-resolution [71]. Some researchers address this by leveraging a priori knowledge from a pre-trained diffusion model [77], compensating for lost texture information in LR images. However, the lack of information constraints coupled with the inherent randomness of the diffusion model contributes to the deviation from fundamental principles.
To address these challenges, we introduce the SSR framework, leveraging depth information to guide blur kernel estimation and subsequently inform the diffusion process [62]. This approach facilitates HR image reconstruction from LR images without compromising the diffusion prior inherent in the generative model. For spatially variant blur kernel estimation, we proposed the Spatially Variant Kernel Refinement (SVKR) module. Within SVKR, Depth-Informed Kernel Estimate Network (DKENet) is devised to utilize the depth information across image locations, and estimate spatially variant blur kernels based on depth information. Introducing depth information enhances the variability of depth information for different objects, aiding in characterizing target object contours. Additionally, SVKR further refine the accuracy of estimated kernel and depth information from blurred, low-resolution images. Moreover, to fully exploit the value of depth information, we have developed the Adaptive Multi-Modal Fusion (AMF) module to merge information from multiple modalities, thereby constraining the diffusion model with spatially variant degradation.
@@
In summary, our main contributions are four-fold: 1) We propose an blind image super-resolution framework known as SSR. SSR employs multiple scales of encoders for information from different modalities using the AMF module, thus generate images with higher fidelity without compromising the diffusion priors. 2) We proposed a Spatially Variant Kernel Refinement (SVKR) module to estimate blur kernels. SVKR iteratively enhance the LR images, thereby refining the precision of blur kernel and depth information estimation. 3) We developed Adaptive Multi-Modal Fusion (AMF) module to merge information from multiple modalities, thereby constraining the diffusion model to generate more authentic SR results. 4) Extensive quantitative and qualitative experiments on representative datasets have verified the superior performance and effectiveness of our method.
@@ 2.1 Non-Blind SR
To enhance non-blind super-resolution, various methods have been developed [34, 78]. SRMD [73] introduced intricate blur kernel assumptions and incorporated them into the network. SFTMD [17] fused blur kernels using SFT layers, while DPSR [74] offered a plug-and-play framework for uniform blur kernels. UDVD [45] used dynamic convolution to address varying degradations, and DPIR [72] employed a HQS to iteratively address data subproblems and prior subproblems. Additionally, zero-shot super-resolution methods have emerged. ZSSR [3] introduced unsupervised CNN super-resolution, and DGDML-SR [13] used depth information for network training. However, these methods face challenges related to foundational assumptions. For example, it is inaccurate for DGML-SR to determine the resolution as high or low solely based on distance. Although these non-blind super-resolution methods have made commendable progress, they often exhibit disappointing performance when applied to complex real-world scenarios.
@@ 2.3 Diffusion Model Based SR
Denoising Diffusion Probability Models (DDPM) [18] have introduced a new generation paradigm, leveraging Markov chains to create stable inverse operations for complex generative processes. Researchers have used DDPM for image super-resolution [16, 23]. SR3 [43] adopts a supervised diffusion model-based approach, coupling degraded images with generated images to facilitate conditional image generation for super-resolution. In contrast, ILVR [15] replaces the low-frequency components in the denoising output with those from the reference image, utilizing an unconditional diffusion model for no-training conditional generation in image super-resolution. CDPMSR [39] improves low-quality images by leveraging pre-existing super-resolution models, providing better conditions for the diffusion model.
Given the intrinsic demand for high fidelity in super-resolution (SR), generated images must maintain a strong consistency with real images. However, these methods employ LR images or enhanced LR images as the diffusion conditions, resulting in a disparity between the images generated in the sampling space of the diffusion model and real images [24]. Our approach introduces depth maps and blur kernels as additional conditioning information, thereby constraining the sampling space, and rendering the generated images more closely aligned with real images.
@@ 3.1 Overview
As previously stated, the application of a spatially variant kernel and depth information can significantly enhance the accuracy of image super-resolution tasks. Pre-trained diffusion models encapsulate an extensive repository of a priori knowledge concerning intricate textures. Drawing inspiration from this premise, we introduce a framework named Adaptive Multi-modal Fusion of Spatially Variant Kernel Refinement with Diffusion Model for Blind Image Super-Resolution (SSR). Through the integration of multi-modal data, we impose constraints on the diffusion model’s sampling space, thereby guiding the diffusion process to produce images that closely resemble reality.
Our initial step involves the iterative estimation of depth information and spatially variant blur kernels from low-resolution (LR) images with DKENet. Specifically, within the iterative estimation process, we first infer the depth map and estimate the spatially variant blur kernels. Subsequently, we combine the blur kernel with a non-blind super-resolution method based on RRDB [56] to estimate a high-resolution image. From this high-resolution image, we once again estimate the depth information. Iterating this process ensures that the estimation of depth information and the blurred kernel mutually reinforce each other, thereby enhancing the precision of the guidance information. Meanwhile, inspired by StableDiffusion, we encode LR images into a latent space using a VAEGAN encoder, and this encoded information guides the diffusion model. These process produce three modolities of LR image, Depth map and blur kernel. Subsequently, information from three different modalities is encoded using encoders of varying scales and is aligned through an Adaptive Multi-Modal Fusion (AMF) module. This aligned information is employed to guide the diffusion process. Through AMF, the aligned features can effectively fine-tune the diffusion model, reducing the number of parameters and the complexity of training. The overall architectural diagram is depicted in Figure 3.
@@ 3.2 Spatially Variant Kernel Refinement
Depth-Informed Kernel Estimate Network In the realm of blind super-resolution, the high-resolution image $x$ and the kernel $k$ in Equation (1) are unknown. We need to estimate $k$ and then use it as prior knowledge to reconstruct $x$. For an image $x\in\mathcal{R}^{H\times W}$, we undertake the estimation of a kernel $k\in\mathcal{R}^{hw\times H\times W}$, where $h$ and $w$ respectively denote the height and width of the kernel. Each component of $k$ corresponds to an individual pixel within $x$.
We utilize Depth-Informed Kernel Estimate Network (DKENet) to estimate the Depth-Informed Kernel. The Depth-Informed Kernel is a spatially variant blur kernel that incorporates depth information of the image during its estimation process. DKENet is structured into two distinct branches, each addressing a separate input modality: one branch processes LR images, while the other extracts features from depth maps. In the LR image processing branch, the network’s operation commences with a standard $3\times 3$ convolutional layer, purposed for preliminary feature extraction. Subsequently, the input features are channeled into a Depth-Aware Block (DABlock), enhancing the network’s responsiveness to pivotal features.
@@
Following this, the feature map undergoes a downsampling operation to reduce its spatial resolution, and then passes through another DABlock for further feature refinement. Subsequent upsampling increases the resolution of the feature map, ensuring the density of information. After another pass through a DABlock, the features proceed through an additional convolutional layer and culminate in a Softmax layer and a upscale-layer. Regarding the depth map branch, it undergoes a feature extraction process similar to that of the LR image branch, ultimately converging with the LR image branch through an SFT layer.
Spatially Variant Kernel Refinement The acquisition of depth information often requires specialized equipment like Time-of-Flight cameras and laser rangefinders [37, 41], which are not always practical for image super-resolution applications [46, 58]. Thus, we use a monocular depth estimation algorithm to derive depth information from low-resolution images. It’s crucial to note that image resolution significantly influences the accuracy of depth estimation; acquiring depth from a blurry, low-resolution image is inherently imprecise [40, 48]. To overcome this, we have devised Spatially Variant Kernel Refinement (SVKR) to augment low-resolution images, subsequently deriving depth information from the resulting images. This intricate process synergizes depth information to facilitate the estimation of the blur kernel.
@@
Initialization: $t=0$, $I_{0}=LR$; $D^{0}=DE(I^{0})$; $K^{0}=KE(LR,D^{0})$; while $t<T$ do
       Estimate depth map: $D^{t+1}\leftarrow DE(I^{t})$; Estimate kernel: $K^{t+1}\leftarrow KE(LR,D^{t+1})$; $K^{t+1}\leftarrow\gamma K^{t+1}+(1-\gamma)K^{t}$;
@@
Algorithm 1 shows the process of SVKR. $DE$ is the current state-of-the-art monocular estimation algorithm, while $KE$ assumes the role of a spatially varying kernel estimation algorithm, and $NBSR$ stands as a lightweight, RRDB-based non-blind super-resolution method. In the course of these iterations, $DE$ estimates the depth map from the most recent SR image $I^{t}$, $KE$ harnesses the most recent depth map $D^{t}$ to estimate the blur kernel from the LR image, and $NBSR$ amalgamates the latest blurred kernel $K^{t}$ to enhance the super-resolved image. This intricate coordination effectively combines the estimations of depth maps and blur kernels, mutually enhancing each other’s precision.
When $t\geq 1$, $K$ is updated using Equation (2), with $\gamma$ serving as the hyperparameter that governs the update rate. This approach allows for the retention of the influence of historical information while updating $K$, rendering it persistent in the direction of updates and mitigating the instability of results caused by oscillations during the updating process. Figure 4 demonstrates the effect of iterative enhancement of depth information and degradation information by SVKR.
@@
3.3 Adaptive Multi-Modal Fusion
Due to the distinct structural and characteristic attributes exhibited by data from various modalities, we have carefully devised encoders of diverse scales and architectures. We use a half UNet as the encoder to encode LR images. This encoder comprises the Input Block (IB), the Middle Block (MB), and the Feature Translate Block (FTB). Among the three modalities of information, namely LR images, depth maps, and blur kernels, LR images serve as the primary modality, containing the most abundant repository of information. Hence, we have tailored the largest-scale encoder to accommodate this primary modality. Conversely, depth maps and blur kernels carry a relatively limited amount of informational conten, thus warranting the implementation of smaller-scale encoders. In order to amalgamate the information from these three modalities and employ it for guiding the diffusion process, we have designed the adaptive multi-modal Fusion (AMF) module, as depicted in Equation (3), where $i\in\{1,2,3\}$.
@@ 3.3 Adaptive Multi-Modal Fusion
where $f_{1}$, $f_{2}$, and $f_{3}$ respectively denote the features obtained through the encoding of the blurred kernel, depth map, and LR image. The symbol $BN(\cdot)$ signifies batch normalization, while $M_{\alpha}$ and $M_{\beta}$ denote MLPs. Specifically, $M_{\alpha}$ serves to rescale features along the channels, while $M_{\beta}$ applies an offset to features, ensuring the dynamic adjustment of semantic information for achieving alignment across three modalities in the semantic space. For the LR image denoted as $x$, the objective of the SR diffusion model is to acquire the posterior distribution $p_{\theta}(z_{t}-1|x,z_{t})$ at a given time step $t$, while satisfying the initial condition $z_{0}~{}q(z|x)$, where $z$ represents the corresponding HR image. In accordance with the AMF framework, this objective may be articulated as Equation (4).
@@
Where $u$ serves as the output of the diffusion model, $l=\mathbb{E}_{t,z_{0},\epsilon}[||\epsilon-\epsilon_{\theta}(\sqrt{\hat{\alpha}_{t}}z_{0}+\epsilon\sqrt{1-\hat{\alpha}_{t}},t)||_{2}^{2}]$ represents the original loss function constraining the diffusion model, $\lambda_{i}$ corresponds to the scaling factors, and $C_{i}$ denotes the constraint process applied to feature $f_{i}$.

### v2

@@ Abstract
Pre-trained diffusion models utilized for image generation encapsulate a substantial reservoir of a priori knowledge pertaining to intricate textures. Harnessing the potential of leveraging this a priori knowledge in the context of image super-resolution presents a compelling avenue. Nonetheless, prevailing diffusion-based methodologies presently overlook the constraints imposed by degradation information on the diffusion process. Furthermore, these methods fail to consider the spatial variability inherent in the estimated blur kernel, stemming from factors such as motion jitter and out-of-focus elements in open-environment scenarios. This oversight results in a notable deviation of the image super-resolution effect from fundamental realities. To address these concerns, we introduce a framework known as Adaptive Multi-modal Fusion of Spatially Variant Kernel Refinement with Diffusion Model for Blind Image Super-Resolution (SSR). Within the SSR framework, we propose a Spatially Variant Kernel Refinement (SVKR) module. SVKR estimates a Depth-Informed Kernel, which takes the depth information into account and is spatially variant. Additionally, SVKR enhance the accuracy of depth information acquired from LR images, allowing for mutual enhancement between the depth map and blur kernel estimates. Finally, we introduce the Adaptive Multi-Modal Fusion (AMF) module to align the information from three modalities: low-resolution images, depth maps, and blur kernels. This alignment can constrain the diffusion model to generate more authentic SR results.
@@ 1 Introduction
Image super-resolution is a foundational pursuit within the domain of computer vision, focusing on the reconstruction of high-resolution images from their low-resolution counterparts to significantly augment pixel density [7, 6]. This intricate task involves the inference and restoration of high-frequency details that are often obscured within the constraints of limited pixel resolutions. Mathematically, the regression process from a high-resolution image (HR) to a low-resolution image (LR) is modeled as follows:
@@
where $k$ symbolizes the blur kernel, $\otimes$ denotes the convolution operation, $\downarrow_{s}$ signifies a downsampling by a factor of $s$, $n$ represents noise. Over the past decade, methods grounded in CNN or transformers have made remarkable strides [11, 13, 34, 70]. Traditional super-resolution methods assume this degradation is characterized by a specific blur kernel, such as a bicubic kernel [35]. However, when a well-trained SR model receives LR images away from its training distribution, it often suffers significant performance degradation [26]. To tackle this problem, blind super-resolution tasks aim at image super-resolution with unknown degradation processes [30].
As depicted in Figure 2, spatially invariant estimated blur kernels, incorporating a priori degradation information, are employed in the super-resolution process to enhance performance and achieve a finite solution space for the LR-HR image pair [5, 8]. In real-world imaging, factors limiting image resolution extend beyond specific blur kernels to encompass motion blur, depth variation in imaging objects, and optical characteristics of imaging devices [60, 61, 19, 65]. The use of a spatially invariant blur kernel becomes inadequate in accurately describing this multifaceted process, yielding unsatisfactory results post super-resolution [64]. Some researchers address this by leveraging a priori knowledge from a pre-trained diffusion model [71], compensating for lost texture information in LR images. However, the lack of information constraints coupled with the inherent randomness of the diffusion model contributes to the deviation from fundamental principles.
@@
To address these challenges, we introduce the SSR framework, leveraging depth information to guide blur kernel estimation and subsequently inform the diffusion process [58, 62]. This approach facilitates HR image reconstruction from LR images without compromising the diffusion prior inherent in the generative model. For spatially variant blur kernel estimation, we proposed the Spatially Variant Kernel Refinement (SVKR) module. Within SVKR, Depth-Informed Kernel Estimate Network (DKENet) is devised to utilize the depth information across image locations, and estimate spatially variant blur kernels based on depth information. Introducing depth information enhances the variability of depth information for different objects, aiding in characterizing target object contours. Additionally, SVKR further refines the accuracy of estimated kernel and depth information from blurred, low-resolution images. Moreover, to fully exploit the value of depth information, we have developed the Adaptive Multi-Modal Fusion (AMF) module to merge information from multiple modalities, thereby constraining the diffusion model with spatially variant degradation.
In summary, our main contributions are four-fold: 1) We propose a blind image super-resolution framework known as SSR. SSR employs multiple scales of encoders for information from different modalities using the AMF module, thus generates images with higher fidelity without compromising the diffusion priors. 2) We proposed a Spatially Variant Kernel Refinement (SVKR) module to estimate blur kernels. SVKR iteratively enhance the LR images, thereby refining the precision of blur kernel and depth information estimation. 3) We developed Adaptive Multi-Modal Fusion (AMF) module to merge information from multiple modalities, thereby constraining the diffusion model to generate more authentic SR results. 4) Extensive quantitative and qualitative experiments on representative datasets have verified the superior performance and effectiveness of our method.
@@ 2.1 Non-Blind SR
To enhance non-blind super-resolution, various methods have been developed [32, 22, 72, 29]. SRMD [67] introduced intricate blur kernel assumptions and incorporated them into the network. SFTMD [16] fused blur kernels using SFT layers, while DPSR [68] offered a plug-and-play framework for uniform blur kernels. UDVD [46] used dynamic convolution to address varying degradations, and DPIR [66] employed a HQS to iteratively address data subproblems and prior subproblems. Additionally, zero-shot super-resolution methods have emerged. ZSSR [3] introduced unsupervised CNN super-resolution, and DGDML-SR [12] used depth information for network training. However, these methods face challenges related to foundational assumptions. For example, it is inaccurate for DGML-SR to determine the resolution as high or low solely based on distance. Although these non-blind super-resolution methods have made commendable progress, they often exhibit disappointing performance when applied to complex real-world scenarios.
@@ 2.3 Diffusion Model Based SR
Denoising Diffusion Probability Models (DDPM) [17] have introduced a new generation paradigm, leveraging Markov chains to create stable inverse operations for complex generative processes. Researchers have used DDPM for image SR [15, 24]. SR3 [45] adopts a supervised diffusion model-based approach, coupling degraded images with generated images to facilitate conditional image generation for SR. In contrast, ILVR [14] replaces the low-frequency components in the denoising output with those from the reference image, utilizing an unconditional diffusion model for no-training conditional generation in image SR. CDPMSR [40] improves low-quality images by leveraging pre-existing SR models, providing better conditions for the diffusion model.
Given the intrinsic demand for high fidelity in SR, generated images must maintain a strong consistency with real images. However, these methods employ LR images or enhanced LR images as the diffusion conditions, resulting in a disparity between the images generated in the sampling space of the diffusion model and real images [25]. Our approach introduces depth maps and blur kernels as additional conditioning information, thereby constraining the sampling space, and rendering the generated images more closely aligned with real images.
@@ 3.1 Overview
As previously stated, the application of a spatially variant kernel and depth information can significantly enhance the accuracy of image SR tasks. With the development of deep learning, many network architectures have emerged, including CNN [48, 51], Transformer [49, 37], and diffusion model [17, 36, 15]. Pre-trained diffusion models encapsulate an extensive repository of a priori knowledge concerning intricate textures. Drawing inspiration from this premise, we introduce a framework named Adaptive Multi-modal Fusion of Spatially Variant Kernel Refinement with Diffusion Model for Blind Image Super-Resolution (SSR). Through the integration of multi-modal data, we impose constraints on the diffusion model’s sampling space, thereby guiding the diffusion process to produce images that closely resemble reality.
Our initial step involves the iterative estimation of depth information and spatially variant blur kernels from low-resolution images (LR) with DKENet. Specifically, within the iterative estimation process, we first infer the depth map and estimate the spatially variant blur kernels. Subsequently, we combine the blur kernel with a non-blind SR method based on RRDB [55] to estimate a high-resolution image. From this high-resolution image, we once again estimate the depth information. Iterating this process ensures that the estimation of depth information and the blurred kernel mutually reinforce each other, thereby enhancing the precision of the guidance information. Meanwhile, inspired by StableDiffusion [44], we encode LR images into a latent space using a VAEGAN encoder [23], and this encoded information guides the diffusion model. These process produce three modalities of LR image, Depth map and blur kernel. Subsequently, information from three different modalities is encoded using encoders of varying scales and is aligned through an Adaptive Multi-Modal Fusion (AMF) module. This aligned information is employed to guide the diffusion process. Through AMF, the aligned features can effectively fine-tune the diffusion model, reducing the number of parameters and the complexity of training. The overall architectural diagram is depicted in Figure 3.
@@ 3.2 Spatially Variant Kernel Refinement
Depth-Informed Kernel Estimate Network. In the realm of blind SR, the high-resolution image $x$ and the kernel $k$ in Equation (1) are unknown. We need to estimate $k$ and then use it as prior knowledge to reconstruct $x$. For an image $x\in\mathcal{R}^{H\times W}$, we undertake the estimation of a kernel $k\in\mathcal{R}^{hw\times H\times W}$, where $h$ and $w$ respectively denote the height and width of the kernel. Each component of $k$ corresponds to an individual pixel within $x$.
We utilize Depth-Informed Kernel Estimate Network to estimate the Depth-Informed Kernel. The Depth-Informed Kernel is a spatially variant blur kernel that incorporates depth information of the image during its estimation process. DKENet is structured into two distinct branches, each addressing a separate input modality: one branch processes LR images, while the other extracts features from depth maps. In the LR image processing branch, the network’s operation commences with a standard $3\times 3$ convolutional layer, purposed for preliminary feature extraction. Subsequently, the input features are channeled into a Depth-Aware Block, enhancing the network’s responsiveness to pivotal features.
@@
Following this, the feature map undergoes a downsampling operation to reduce its spatial resolution, and then passes through another DABlock for further feature refinement. Subsequent upsampling increases the resolution of the feature map, ensuring the density of information. After another pass through a DABlock, the features proceed through a convolutional layer and culminate in a Softmax layer and a upscale-layer. Regarding the depth map branch, it undergoes a feature extraction process similar to that of the LR image branch, ultimately converging with the LR image branch through an SFT layer.
Spatially Variant Kernel Refinement. The acquisition of depth information often requires specialized equipment like Time-of-Flight cameras and laser rangefinders [38, 42], which are not always practical for image SR applications [47, 56]. Thus, we use a monocular depth estimation algorithm to derive depth information from low-resolution images. It’s crucial to note that image resolution significantly influences the accuracy of depth estimation; acquiring depth from a blurry, low-resolution image is inherently imprecise [41, 50]. To overcome this, we have devised Spatially Variant Kernel Refinement (SVKR) to augment low-resolution images, subsequently deriving depth information from the resulting images. This intricate process synergizes depth information to facilitate the estimation of the blur kernel.
@@
Initialization: $t=0$, $I_{0}=LR$; $D^{0}=DE(I^{0})$; $K^{0}=KE(LR,D^{0})$;
while $t<T$ do
       Estimate depth map: $D^{t+1}\leftarrow DE(I^{t})$;
      Estimate kernel: $K^{t+1}\leftarrow KE(LR,D^{t+1})$; $K^{t+1}\leftarrow\gamma K^{t+1}+(1-\gamma)K^{t}$;
@@
Algorithm 1 shows the process of SVKR. $DE$ is the current state-of-the-art monocular estimation algorithm, while $KE$ assumes the role of a spatially varying kernel estimation algorithm, and $NBSR$ stands as a lightweight, RRDB-based non-blind SR method. In the course of these iterations, $DE$ estimates the depth map from the most recent SR image $I^{t}$, $KE$ harnesses the most recent depth map $D^{t}$ to estimate the blur kernel from the LR image, and $NBSR$ amalgamates the latest blurred kernel $K^{t}$ to enhance the super-resolved image. This intricate coordination effectively combines the estimations of depth maps and blur kernels, mutually enhancing each other’s precision.
When $t\geq 1$, $K$ is updated using Equation (2), where $\gamma$ is the hyperparameter that governs the update rate. This approach allows for the retention of the influence of historical information while updating $K$, rendering it persistent in the direction of updates and mitigating the instability of results caused by oscillations during the updating process. Figure 4 demonstrates the effect of iterative enhancement of depth information and degradation information by SVKR.
@@
3.3 Adaptive Multi-Modal Fusion.
Due to the distinct structural and characteristic attributes exhibited by data from various modalities, we have carefully devised encoders of diverse scales and architectures. We use a half UNet as the encoder to encode LR images. This encoder comprises the Input Block (IB), the Middle Block (MB), and the Feature Translate Block (FTB). Among the three modalities of information, namely LR images, depth maps, and blur kernels, LR images serve as the primary modality, containing the most abundant repository of information. Hence, we have tailored the largest-scale encoder to accommodate this primary modality. Conversely, depth maps and blur kernels carry a relatively limited amount of informational content, thus warranting the implementation of smaller-scale encoders. In order to amalgamate the information from these three modalities and employ it for guiding the diffusion process, we have designed the adaptive multi-modal Fusion (AMF) module, as depicted in Equation (3), where $i\in\{1,2,3\}$.
@@
where $f_{1}$, $f_{2}$, and $f_{3}$ respectively denote the features obtained through the encoding of the blurred kernel, depth map, and LR image. The symbol $BN(\cdot)$ signifies batch normalization, while $M_{\alpha}$ and $M_{\beta}$ denote MLPs. Specifically, $M_{\alpha}$ serves to rescale features along the channels, while $M_{\beta}$ applies an offset to features, ensuring the dynamic adjustment of semantic information for achieving alignment across three modalities in the semantic space. For the LR image denoted as $x$, the objective of the SR diffusion model is to acquire the posterior distribution $p_{\theta}(z_{t}-1|x,z_{t})$ at a given time step $t$, while satisfying the initial condition $z_{0}\sim q(z|x)$, where $z$ represents the corresponding HR image. In accordance with the AMF framework, this objective may be articulated as Equation (4).
@@
Where $u$ serves as the output of the diffusion model, $l=\mathbb{E}_{t,z_{0},\epsilon}[||\epsilon-\epsilon_{\theta}(\sqrt{\hat{\alpha}_{t}}z_{0}+\epsilon\sqrt{1-\hat{\alpha}_{t}},t)||_{2}^{2}]$ represents the original loss function constraining the diffusion model, $\lambda_{i}$ corresponds to the scaling factors, and $C_{i}$ denotes the L1 regularization applied to feature $f_{i}$. We used L1 regularization because $f_{i}$ are features from different modalities, and sparsifying them helps the network learn complementary information.